OpenStudy (anonymous):

a vehicle with a mass of 800kg experience a friction force of 180N and travels at a constant speed of 54km/h. 1.Calculate the power required from the engine to travel at the constant speed. 2.The work done to accelerate the vehicle from 54km/h with an acceleration of 2 m/s for 5 seconds

OpenStudy (theeric):

1. Power is work per time. Work is force times distance in the direction of force. P= fd/t = f v. I don't have anymore time, but I saw you're on so I decided to give you some information!

OpenStudy (anonymous):

@LinuxMint still need help here?

OpenStudy (anonymous):

hi ive done the following, im 90% sure im on the right track although it took me a while to actually understand it. Power = force x speed, the force is 180N the speed is how fast the car is travelling in metres per second Power = 180 x 54000 / 3600 (the answer is in watts) Work done is force x distance. Force to accelerate vehicle use F = ma = 800 x 2 + 180 (i need to add in the friction of 180N as this must be overcome) F = 1780N. Using s = ut + 1/2 at^2 you can work out how far it travels during this acceleration (just plugged the numbers in!) u =54, a = 2, t = 5 And then multiply it by the force I worked out. what do you guys think?

OpenStudy (anonymous):

Second part is not OK. The mass is already moving at a constant 54Km/h (because the foce of 180N is applied to compensate friction). The work done to accelerate at 2m/sec^2 for 5 secs is additional to the 180N, then you do not have to add 180N (they are already there)

OpenStudy (anonymous):

The additional work contributes to increase Kinetic energy Initial velocity = Vo(54 Km/h) Final Velocity =Vo+at=Vo+10 Increment of kinetic energy = \[\Delta E_k=\frac{ 1 }{ 2 }m(v_0+10)^2-\frac{ 1 }{ 2 }mv_0^2\]

OpenStudy (anonymous):

I get 160 KJoules The formula is coming from: \[dE=F·ds=m·a·ds=m·\frac{ dv }{ dt }ds=m \left( \frac{ ds }{ dt }\right)dv=m·v·dv\]Then Work done between x1 and x2 is:\[W=E=\int\limits_{x_0}^{x_1}Fds=\int\limits_{v_0}^{v_0+10}m·v·dv=\frac{ 1 }{ 2 }m \left[ v^2 \right]_{v_0}^{v_0+10}\]

OpenStudy (anonymous):

I mean work done between xo and x1

OpenStudy (anonymous):

Another way to do it: Distance moved in 5 seconds \[d=v_0t+\frac{ 1 }{ 2 }a·t^2=15·5+\frac{ 1 }{ 2 }·2·25=100\] The additional force is: \[F=m·a=800·2=1600\]Additional work\[W=F·d=1600·100=160,000 J=160KJ\]

OpenStudy (anonymous):

i believe you are correct regarding the issue of friction, i actually understand now how to see the effect of forces upon a vehicle with a constant speed. i think the keyword here was CONSTANT which i missed totally. i see when it comes to calculations its better to simplify before tackling a problem regarding v=54km/h=15m/s. this is my second week studying engineering science. i only started studying after 14 years of laziness. i guess no time like the present. thanks for helping me. much appreciated

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

hi , regarding your explanation of friction effect in the calculation of work done, im thinking about this question and i wonder wouldn't i need to overcome friction just to increase velocity because essentially i am increasing velocity due to my acceleration after i have achieved constant speed?

OpenStudy (anonymous):

Absolutely not. Friction was already overcome, otherwise the mass would not move at a constant speed in the beginning. If the mass is moving at constant speed that means there is no acceleration and therefore no net force on it. If you now apply a force (additional) the effect of this force will be to increase kinetic energy (velocity) Read carefully the problem statement:"The work done to accelerate the vehicle from 54km/h with an acceleration of 2 m/s for 5 seconds" If the 180 Newton force was used to accelarate, then it would not be balancing friction and the initial velocity would not be 54 km/h and of course would not be constant (the mass would be decelerating due to friction)

OpenStudy (anonymous):

There are two energies here: The force used to keep velocity constant (balancing friction) and the force to increase velocity from 15m/sec up to 25 m/sec. It is the work done by this force what we are looking for

OpenStudy (anonymous):

thanks for the help much appreciated , i now understand the question completely and your explanation makes a lot of sense to me . just one question : do you have a website to recommend to study moments, beams, shearing force force diagrams etc. my textbook sucks

OpenStudy (anonymous):

Which level are you studying? Is it high school or college?

OpenStudy (anonymous):

first year college but had studied a couple of years back so i'm quite rusted in terms of learning

OpenStudy (anonymous):

well, I am afraid I studied Physics many years ago and by that time internet was not even invented. What I usually do is refresh things with my books (probably some of them out of print) and google. What is the book you are using?

OpenStudy (anonymous):

engineering science by MJJ van Rensburg

OpenStudy (anonymous):

and is covering only Physics?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im studying part time , so im doing my course per module.

OpenStudy (anonymous):

im doing civil engineering course

OpenStudy (anonymous):

I have always liked the books by Paul Tipler. Good books for a general course in Physics (expensive, though) Also Alonso-Finn are really impressive, advanced (and tough) for Engineering courses (expensive too) Halliday-Resnick is a good option (some people dont like it, I do) "An Introduction to Mechanics" by Kleppner. I do not have it but always heard a very good critique though it is a bit old fashioned format (was used in MIT)

OpenStudy (anonymous):

my current work is in waste water management, and i really enjoy it , pretty exciting stuff. im in the process of advancing my career hence civil . i will check out some of the books recommended. ive also heard good critique about kleppner